# Quantum critical theories in a periodic potential: strange metallic thermoelectric and magneto-transport

### v1

@@ 6 Conclusion and Outlook
Quantum critical systems can give rise to idiosyncratic transport signatures. This study of DC and AC thermoelectric and magneto-transport in 2D quantum critical theories with strong translational symmetry breaking due to a spatially varying chemical potential lattice with zero average exhibits this again. The most remarkable aspect is that the transport is never Drude-like in the sense that there is never a single longest-lived (momentum) relaxation mode controlling transport. In a 1D periodic potential electrical transport is purely quantum critical, whereas thermal transport is governed by two propagating modes originating in a collision of the would-be Drude mode with the Umklapp charge-diffusion mode. In a 2D lattice, on the other hand, the low-energy physics resembles that of Effective Medium Theory: currents follow meandering? spatial paths of least resistance in a heterogeneous system built out of local regions that resemble alternating particle- and hole filled states. This results in new weak diffusive modes that are not the global momentum-relaxation mode, but probably better seen as remnants of the local Drude modes from of the local charge-filled regions. This resemblance to results from Effective Medium Theory is also present in magneto-transport, where the longitudinal magnetoresistance shows a non-analytic unsaturating scaling behavior at large $B$ that is approximately linear.
@@
The context of this physics is familiar from studies of charge neutral graphene on various substrates with a periodicity larger than the intrinsic lattice. These studies have also shown the connection with Effective Medium Theory phenomenology cheianovRandomResistorNetwork2007a; rossiEffectiveMediumTheory2009; fernandesEffectiveMediumModel2023. Our findings that one can have to the eye Drude-like transport co-existing with unsaturating $B$-linear-like magnetoresistance may be also relevant to the strange metal phase of high-$T_{c}$ cuprates. It provides a potential answer to the two-time scale puzzle needed to explain differing phenomenology of Hall and longitudinal transport, which has stubbornly eluded resolution since it was pointed out by Anderson. A purely charge-neutral system as studied here will however not apply; for one, the Hall effect vanishes. More fundamentally, theoretical studies in SYK models have shown that one must have an IR fixed point determined by $SL(2,R)$ time-reparametrization invariance to explain local quantum criticality. In terms of its holographic dual, the $T=0$ near horizon geometry must be AdS2, whereas charge-neutral systems have an AdS4 geometry. Once a finite density is added to the problem, the question to answer is whether a low-temperature regime with AdS2 physics might still give Effective Medium Theory multi-time-scale signatures, as long as $T/G\gtrsim 1$. [^13] One can already deduce that for the characteristic temperature of a cuprate strange metal $T=100K$, the regime $T/G=1$ corresponds roughly to spatial variations on the order of $500\,\mathrm{nm}$ assuming the relevant speed is the Fermi velocity $v_{\text{F}}\sim 5\cdot 10^{6}\,\mathrm{m/s}$. The relevant modulation must therefore occur on scales far larger than the atomic lattice. A very recent SYK-based computation valentinisSuperlinearHallAngle2025 indeed finds different time-scales in ordinary and Hall transport, due to an effective increase of the Hall coefficient $R_{H}$ with $T$. As $R_{H}=B/n$ perturbatively, this naively implies zero-density at $T=0$, suggestively indicating that the fundamental physics is similar. We shall address this in a future study.
@@ Appendix A Numerical method
The equations of motion were obtained in Mathematica with the help of the xAct package brizuela_xpert_2009. To solve the resulting linear/non-linear equations of motion, we have used the PETSc library in c petsc-efficient; petsc-user-ref. For a guide on implementation of the latter we refer the reader to Ref. bueler_petsc_2021.
@@
which clusters points near the boundary and the horizon. In the transverse $xy$-directions, we used 8th/4th order finite differences on uniform grids for 1D/2D lattices. This leads to additional sparsity in the matrix discretizing the problem, and most importantly, allows for efficient partitioning of the matrix on multiple processors. PETSc is built for finite differencing, so to implement the radial pseudospectral derivatives we force PETSc not to split the domain along the $z$-direction across the processors. [^14] This is achieved by first lying to PETSc’s DMDA routine that the stencil width is that of the finite difference stencil and then explicitly telling PETSc how many non-zero entries per row there are (otherwise a lot of time is lost on memory allocation).
To solve the non-linear equations for the backgrounds we used a GMRES non-linear solver, as the matrices are in general not symmetric due to the boundary conditions in the radial coordinate. To solve the linear equations (at each step of the non-linear algorithm, or for the linear equations of the perturbations), we used Schwarz domain decomposition implemented via an ASM preconditioner, distributing the calculation over 4-10 processors. As a sub-preconditioner on each block, we used ILU or LU factorization with varying levels of overlap, depending on the numerical difficulty. The linear equations for the perturbations often require full LU factorization with 2 points of overlap, whereas it generally suffices with ILU(0) with 1 point of overlap for the non-linear background equations, expect for the very strongest lattices. The Stokes equations are easily solved with exact LU factorization on a single process. Note that since the Stokes equations 12 only depend on the derivatives of $p$ and $w$, one needs to fix a value for these functions at some point in the domain in order to remove a pair of zero modes of the matrix discretizing the equations.
@@ A.1 Numerical convergence
The numerical difficulty increases as the temperature/lattice is reduced/increased. Throughout our calculations, we use units set by the horizon radius, which we fix to unity. In these units, the temperature is constant (for $B=0$), so lowering $T/G$ corresponds to increasing the lattice vector $G$. Maintaining a large ratio $\delta\mu/G$ therefore becomes increasingly challenging at low temperatures, since this requires $\delta\mu$ (in units of the horizon radius) to be correspondingly large.

### v2

@@ 6 Conclusion and outlook
Quantum critical systems can give rise to idiosyncratic transport signatures. This study of DC and AC thermoelectric and magneto-transport in 2D quantum critical theories with strong translational symmetry breaking due to a spatially varying chemical potential lattice with zero average exhibits this again. The most remarkable aspect is that the transport is never Drude-like in the sense that there is never a single longest-lived (momentum) relaxation mode controlling transport. In a 1D periodic potential electrical transport is purely quantum critical, whereas thermal transport is governed by two propagating modes originating in a collision of the would-be Drude mode with the Umklapp charge-diffusion mode. In a 2D lattice, on the other hand, the low-energy physics resembles that of Effective Medium Theory: currents follow meandering spatial paths of least resistance in a heterogeneous system built out of local regions that resemble alternating particle-filled and hole-filled states. This results in new weak diffusive modes that are not the global momentum-relaxation mode, but probably better seen as remnants of the local Drude modes from of the local charge-filled regions. This resemblance to results from Effective Medium Theory is also present in magneto-transport, where the longitudinal magnetoresistance shows a non-analytic unsaturating scaling behavior at large $B$ that is approximately linear.
@@
The context of this physics is familiar from studies of charge neutral graphene on various substrates with a periodicity larger than the intrinsic lattice. These studies have also shown the connection with Effective Medium Theory phenomenology Cheianov et al. (2007); Rossi et al. (2009); Fernandes (2023). Our findings that one can have to the eye Drude-like transport co-existing with unsaturating $B$-linear-like magnetoresistance may be also relevant to the strange metal phase of high-$T_{c}$ cuprates. It provides a potential answer to the two-time scale puzzle needed to explain differing phenomenology of Hall and longitudinal transport, which has stubbornly eluded resolution since it was pointed out by Anderson. A purely charge-neutral system as studied here will however not apply; for one, the Hall effect vanishes. More fundamentally, theoretical studies in SYK models have shown that one must have an IR fixed point determined by $SL(2,R)$ time-reparametrization invariance to explain local quantum criticality. In terms of its holographic dual, the $T=0$ near horizon geometry must be AdS2, whereas charge-neutral systems have an AdS4 geometry. Once a finite density is added to the problem, the question to answer is whether a low-temperature regime with AdS2 physics might still give Effective Medium Theory multi-time-scale signatures, as long as $T/G\gtrsim 1$. [^14] One can already deduce that for the characteristic temperature of a cuprate strange metal $T=100K$, the regime $T/G=1$ corresponds roughly to spatial variations on the order of $500\,\mathrm{nm}$ assuming the relevant speed is the Fermi velocity $v_{\text{F}}\sim 5\cdot 10^{6}\,\mathrm{m/s}$. The relevant modulation must therefore occur on scales far larger than the atomic lattice. A very recent SYK-based computation Valentinis et al. (2025) indeed finds different time-scales in ordinary and Hall transport, due to an effective increase of the Hall coefficient $R_{H}$ with $T$ — a feature that also follows from a phenomenological match of cuprate DC transport to hydrodynamics Amoretti et al. (2020). As $R_{H}=B/n$ perturbatively, this naively implies zero-density at $T=0$, suggestively indicating that the fundamental physics is similar.
@@ Appendix A Numerical method
The equations of motion were obtained in Mathematica with the help of the xAct package Brizuela et al. (2009). To solve the resulting linear/non-linear equations of motion, we have used the PETSc library in c Balay et al. (1997, 2024). For a guide on implementation of the latter we refer the reader to Ref. Bueler (2021).
@@
which clusters points near the boundary and the horizon. In the transverse $xy$-directions, we used 8th/4th order finite differences on uniform grids for 1D/2D lattices. [^15] In order to resolve the sharp Umklapp sound peak at large $T/G$ in the 2D AC response, we must increase the finite difference order in the $z$-direction to 8. This leads to additional sparsity in the matrix discretizing the problem, and most importantly, allows for efficient partitioning of the matrix on multiple processors. PETSc is built for finite differencing, so to implement the radial pseudospectral derivatives we force PETSc not to split the domain along the $z$-direction across the processors. [^16] This is achieved by first lying to PETSc’s DMDA routine that the stencil width is that of the finite difference stencil and then explicitly telling PETSc how many non-zero entries per row there are (otherwise a lot of time is lost on memory allocation).
To solve the non-linear equations for the backgrounds we used a GMRES non-linear solver, as the matrices are in general not symmetric due to the boundary conditions in the radial coordinate. To solve the linear equations (at each step of the non-linear algorithm, or for the linear equations of the perturbations), we used Schwarz domain decomposition implemented via an ASM preconditioner, distributing the calculation over 4–10 processors. As a sub-preconditioner on each block, we used ILU or LU factorization with varying levels of overlap, depending on the numerical difficulty. The linear equations for the perturbations often require full LU factorization with 2 points of overlap, whereas it generally suffices with ILU(0) with 1 point of overlap for the non-linear background equations, expect for the very strongest lattices. The Stokes equations are easily solved with exact LU factorization on a single process. Note that since the Stokes equations 12 only depend on the derivatives of $p$ and $w$, one needs to fix a value for these functions at some point in the domain in order to remove a pair of zero modes of the matrix discretizing the equations.
@@ A.1 Numerical convergence
The numerical difficulty increases as the temperature/lattice strength is reduced/increased. Throughout our calculations, we use units set by the horizon radius, which we fix to unity. In these units, the temperature is constant (for $B=0$), so lowering $T/G$ corresponds to increasing the lattice vector $G$. Maintaining a large ratio $\delta\mu/G$ therefore becomes increasingly challenging at low temperatures, since this requires $\delta\mu$ (in units of the horizon radius) to be correspondingly large.